OpenStudy (anonymous):

Identify the maximum and minimum values of the function y = 8 cos x in the interval [-2π, 2π]. Use your understanding of transformations, not your graphing calculator

OpenStudy (harsimran_hs4):

cool this is rather simple ..... can you tell me the maximum and minimum value of cosx in the interval [-2π, 2π] ?

OpenStudy (anonymous):

there are many ways to solve this problem, one way is to sketch the function and spot the max/min, or you may simply use calculus by differentiating the function and setting it to zero

OpenStudy (anonymous):

have a look at the plot, and identify the max mountains and min valleys

OpenStudy (anonymous):

so how should i put it in for the answer

OpenStudy (anonymous):

You know that cosine is bounded: \[-1\le\cos x\le1\] What can you say about 8*cosine? \[?\le8\cos x\le?\]

OpenStudy (anonymous):

rite

OpenStudy (anonymous):

so put it just like that

OpenStudy (anonymous):

Just like what?

OpenStudy (anonymous):

can just only put the answer u gave instead of the other part im confused

OpenStudy (anonymous):

No, I gave you an interval. You're asked for the maximum and minimum values of 8cosx. If you figure out what the two question marks should be, then you have your answer.

OpenStudy (anonymous):

rite

OpenStudy (anonymous):

so it woulld be -2pi< 8 cos x <2pi rite

OpenStudy (anonymous):

No, that's not it. \[-1\le\cos x\le1\\ 8(-1)\le8(\cos x)\le8(1)\\ -8\le8\cos x\le8\] This tells you that the highest value attained by 8cosx is 8 and the lowest value is -8.

OpenStudy (anonymous):

so it would be min= -8 and max= 8 8 cos x

OpenStudy (anonymous):

Yes. The max should be just 8, not "8 8 cosx," though.

OpenStudy (anonymous):

so thats all for this problem

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

you should not get confused between the amplitude and frequencies, or y a-xis and x-axis @a2d...

OpenStudy (anonymous):

if you look as well to the graph, you can see +8 on the y-axis, and -8

OpenStudy (anonymous):

this is another way to visualize max and min

OpenStudy (anonymous):

and this is how your AM radio works in communication

OpenStudy (anonymous):

rite

OpenStudy (anonymous):

go ahead plz

OpenStudy (anonymous):

i sent you the graph at the top as well, but the method that Sith gave you is a good one and also important, if you look at the plot you can see on the y axis you have the max is 8 and the min -8

OpenStudy (anonymous):

@mathsmind @SithsandGiggles I understand how to get 8 and -8. From there, how do we figure out the x value? (x, 8) and (x, -8)